# Gradient is All You Need: Gradient-Based Attention Fusion for Infrared Small Target Detection

### v1

@@ III-C1 Studies of Dilation Rate for GradFormer
ACM[10] 0.40 69.69 82.10 96.82 8.12 63.78 77.91 93.93 5.81
RDIAN[17] 0.90 86.67 93.81 97.98 4.76 64.08 78.12 93.34 5.30
ALCNet[18] 0.37 92.39 96.04 98.83 3.79 65.10 78.88 93.26 7.90
AGPCNet[19] 13.36 86.71 92.88 98.09 3.93 67.54 80.64 94.27 4.76
DNANet[11] 4.69 93.41 96.59 98.94 3.52 65.79 79.23 93.60 5.96
UIU-Net[12] 50.54 94.27 97.04 98.73 1.22 68.06 80.91 92.59 5.85
SCTransNet[20] 11.19 93.38 96.57 97.88 4.73 67.97 80.90 93.26 6.24
Ours 4.46 94.82 97.34 98.94 1.82 68.26 81.11 95.95 4.42
III-C2 Studies of Various Components
We illustrate the effectiveness of each module in Tab. II. The performance of our network improves progressively with the inclusion of each module. GradFormer, which focuses on local difference features, enhances the model’s performance more effectively than GFEM, which primarily relies on global perception. In IRSTD, the capability to learn gradient information is more critical than access to global context.
III-D Comparsion with State-of-the-art (SOTA) Methods
To demonstrate the effectiveness of our proposed model, we compare it with several SOTA methods on two datasets, including SCTransNet [20], UIUNet [12], ACM [10], ALCNet [18], AGPCNet [19], RDIAN [17], and DNANet [11]. As shown in Tab. III, and our method performs excellently. We utilize GaNet with dilation ratios of 1 and 5 as the baseline models for comparative evaluation.
III-D1 Qualitative Result
As shown in Fig. 5, our model preserves the shape of targets with low missed detections (green boxes) and false alarms (yellow boxes) compared to other methods. Our approach excels in learning rich feature representations by seamlessly integrating detailed features with high-level semantics and leveraging a global attention, which enables precise differentiation between targets and backgrounds, effectively mitigating false alarms and reducing missed detections. Despite the challenges posed by small and dim targets in infrared images with complex backgrounds, our network consistently balances segmentation quality with detection accuracy. While methods like UIU-Net may retain target structure, they often suffer from false positives and negatives. Other approaches, such as ALCNet, may exhibit fewer false alarms and missed targets but struggle to maintain target shapes. Through comparisons with other methods, we have highlighted the superior performance of our approach.
@@ III-D2 Quantitative Results
As shown in Tab. III, our method achieves impressive performance, including an $mIoU$ of $95.21\%$, an $F1$ score of $97.51\%$, and a $Pd$ of $99.04\%$, a $Pa$ of $1.06\times 10^{-6}$. With $4.46$M parameters, it is significantly more efficient than other methods. Similarly, on the IRSTD-1k dataset, our method is a leading network. Generally, methods that rely on data with more complex structures tend to perform better. However, our network demonstrates that such models often extract features that are irrelevant to IRSTD. Instead, we focus on utilizing gradients from lower layers and semantics from deeper layers for optimal performance. Additionally, we employ the ROC curve to analyze the performance of various models, as depicted in Fig. 6. While UIU-Net achieves commendable results in $mIoU$, its AUC scores are comparatively weak. In contrast, ALCNet’s performance shows significant divergence, with subpar outcomes in $mIoU$ but relatively strong performance in ROC metrics. The model proposed in this article consistently demonstrates good performance in both ROC and $mIoU$.
Figure 6: ROC curves of varied networks on NUDT-SIRST.
@@ IV Conclusion
In this letter, we propose an IRSTD network that enhances detection performance compared to baseline models and has fewer model parameters. The GradFormer module was proposed to improve the detection performance of the network. Additionally, we propose the GFEM module, which focuses on global perception across the entire feature map and learns relationships between distant pixels, preventing GradFormer from overlooking background information due to its emphasis on local differences. To demonstrate the effectiveness of the proposed network and modules, we conducted extensive experiments that indicate satisfied results and significant practical application potential.

### v2

@@ III-D1 Qualitative Result
ACM[6] 0.40 70.97 82.99 97.67 7.33 63.78 77.91 93.93 5.81
RDIAN[13] 0.90 87.78 93.47 97.88 9.67 64.08 78.12 93.34 5.30
ALCNet[14] 0.37 92.45 96.08 98.94 2.62 65.10 78.88 93.26 7.90
AGPCNet[15] 13.36 85.38 92.11 97.88 5.95 67.54 80.64 94.27 4.76
DNANet[7] 4.69 93.73 96.70 99.25 4.55 65.79 79.23 93.60 5.96
UIU-Net[8] 50.54 94.11 96.96 97.98 0.74 68.06 80.91 92.59 5.85
SCTransNet[16] 11.19 93.83 96.77 98.20 1.56 67.97 80.90 93.26 6.24
Ours 4.46 94.93 97.39 99.04 2.00 68.26 81.11 95.95 4.42
@@ III-D2 Quantitative Results
As shown in the Tab. IV, our method achieves impressive performance, including an $mIoU$ of $94.93\%$, an $F1$ score of $97.39\%$, and a $Pd$ of $99.04\%$, a $Pa$ of $2.00\times 10^{-6}$. Similarly, on the IRSTD-1K dataset, our method is a leading network.
Figure 6: ROC curves of varied networks on NUDT-SIRST datasets.
Furthermore, we used the ROC curve to analyze the performance of various models, as depicted in Fig. 6. Although UIU-Net achieves commendable results in $mIoU$, its AUC scores are comparatively weak. In contrast, ALCNet performance shows a significant divergence, with subpar outcomes in $mIoU$ but relatively strong performance in ROC metrics. The ROC curve corroborates the findings in Tab. IV. The model proposed in this article consistently shows good performance in both ROC and $mIoU$.
@@ IV Conclusion
In this letter, we propose an ISTD network that enhances detection performance. The DATrans was proposed to enhance the local gradient feature extraction and detection performance of the network. Additionally, we propose the GFEM, which focuses on global perception across the entire feature map and learns relationships between distant pixels, preventing DATrans from overlooking background information due to its emphasis on local gradient differences. Then the local and global feature are fused. We conducted extensive experiments indicating satisfied results with fewer parameters. The model is a back-box approach because its capability for dynamic weight adjustments lacks mathematical proof. Moreover, it suffers from false alarms when processing high-noise images. To address these limitations, we aim to improve the model in two aspects. At first, we plan to improve the interpretability of the network. Then, we would like to integrate temporal information into the model to improve its robustness.